PSa 2-11 Calculate Gross Pay for a Piecework Employee

Calculate gross pay for each of the following employees. All are paid an overtime wage rate that is 1.5 times their respective regular wage rates.

NOTE: For simplicity, all calculations throughout this exercise, both intermediate and final, should be rounded to two decimal places at each calculation.

1: Walter Pinkman assembles merchandise and is paid $0.11 for each unit assembled. During the most recent week, he worked 46 hours and assembled 5,628 units.

Gross Pay = $

2: Sidney Darling is a telemarketer, who is paid $0.34 for every telemarketing call he places. During the most recent week, he worked 41 hours and placed 1,642 calls.

Gross Pay = $

3: Pete Brees assembles merchandise and is paid $0.04 for each unit assembled. During the most recent week, he worked 52 hours and assembled 13,284 units.

Gross Pay = $

4: Roy Carter is a telemarketer who is paid $0.30 for every telemarketing phone call he places. During the most recent week, he worked 51 hours and placed 1,536 calls.

Gross Pay = $
Answer #1

1. answer is $659.46

regular earnings for all hours = 5628 units * $0.11 per unit = $619.08

regular hourly rate = 619.08/46 hours = 13.46

overtime rate per hour = 13.46*0.5 = 6.73

overtime earnings = 6 hours * 6.73 = 40.38

total earnings = regular earnings + overtime earnings = 619.08+40.38 = 659.46

2. answer is $565.09

regular earnings for all hours = 1642 calls * $0.34 per call = $558.28

regular hourly rate = 558.28/41 hours = 13.62

overtime rate per hour = 13.62*0.5 = 6.81

overtime earnings = 1 hours * 6.81 = 6.81

total earnings = regular earnings + overtime earnings = 558.28+6.81 = 565.09

3. answer is $592.68

regular earnings for all hours = 13284 units * $0.04 per unit = $531.36

regular hourly rate = 531.36/52 hours = 10.22

overtime rate per hour = 10.22*0.5 = 5.11

overtime earnings = 12 hours * 5.11 = 61.32

total earnings = regular earnings + overtime earnings = 531.36+61.32 = 592.68

4. answer is $510.52

regular earnings for all hours = 1536 calls * $0.30 per call = $460.80

regular hourly rate = 460.80/51 hours = 9.04

overtime rate per hour = 9.04 * 0.5 = 4.52

overtime earnings = 11 hours * 4.52 = 49.72

total earnings = regular earnings + overtime earnings = 460.80+49.72 = 510.52